OpenStudy (anonymous):

Solve the system of equations. 2x – 5y = -3 -x + 2y = 3

OpenStudy (anonymous):

Try substitution.

OpenStudy (anonymous):

solve equation 2 for x x=2y-3 now put this in first equation and solve for y 2(2y-3)-5y=3 4y-6-5y=3 -y=9 y=-9 since x=2y-3 so x=2(-9)-3 x=-18-3=-21 so x=-21 y=-9

OpenStudy (anonymous):

=)) thank youuu

OpenStudy (anonymous):

@coolaidd sorry there was a typo mistake while solving equation for y i missed - on the right side correct solution is solve equation 2 for x x=2y-3 now put this in first equation and solve for y 2(2y-3)-5y=-3 4y-6-5y=-3 -y=3 y=-3 now put in x=2y-3 x=2(-3)-3 x=-6-3 x=-9 x=-9 y=-3